The square of a number is 27 more than 6 times the number
Answer by Boreal(15235)

number is x
x^2-27=6x. That is the square minus 27 equals 6 times it.
Or x^2=6x+27. That works, too.
x^2-6x-27=0. When a square appears, think quadratic equation, and set all the terms on one side.
(x-9)(x+3)=0 factor it, if you can.
The numbers are 9 and -3
Check each
9^2-27=9*6
81-27=54 True.
(-3)^2-27=(-3)(6)
9-27=-18 True